OpenStudy (anonymous):

How long will it take for a quantity of money (A) to triple in value if it is invested at an annual interest rate (r) compounded continuously?

OpenStudy (anonymous):

A= P(1+r)^n we want A =3P

OpenStudy (anonymous):

3=Pe^(r*t)

OpenStudy (anonymous):

ln(3)=ln(e^(r*t)

OpenStudy (anonymous):

so 3P = P (1+r)^n (1+r)^n = 3 n ln ( 1+r) = ln(3) n = ln(3) / [ ln(1+r) ] where r is the rate as a decimal

OpenStudy (anonymous):

So what would be the formula for tripling time?

OpenStudy (anonymous):

yes, e^ you can use decimal

OpenStudy (anonymous):

thats it :|

OpenStudy (anonymous):

n is a time

OpenStudy (anonymous):

if 100=3(100)

OpenStudy (anonymous):

example 300=100e^(r*t)

OpenStudy (anonymous):

p=principal R=percent t+ time

OpenStudy (anonymous):

sorry: t=yime

OpenStudy (anonymous):

I'm confused with what to answer the question I was given with, because my instructor is asking for just a formula, not to evaluate.

OpenStudy (anonymous):

A^3= Pe^(r*t)

OpenStudy (anonymous):

n = ln(3) / [ ln(1+r) ] thats the answer, it is time as a function of rate "Nancy Lam" doesnt know pellet

OpenStudy (anonymous):

depent on you use I use r for interest percent, you can use n

OpenStudy (anonymous):

n is not percent, its a time , geez..you know less than the actual answerer lolz, I get 90+ for engineering maths subjects at uni , who are you going to believe :P

OpenStudy (anonymous):

Thank you, I understand it now!

OpenStudy (anonymous):

you can use any leter

OpenStudy (anonymous):

And thank you Nancy, I appreciate your help!

OpenStudy (anonymous):

Ok, bye